From the information below concerning an employed Kenyan on a pension scheme, calculate the gross pay, chargeable pay, tax charged and PAYE for the individual (Clearly explain and show your workings)


Description AMOUNT


Basic Pay 163,562


ALLOWANCE


Housing Amount 139,675


Commuting Allowance 10,000


Gross Pay ???


TAX CALCULATIONS


Defined Contributions 200


Provident Fund 16,356.2


Taxable Pay ???


Tax Charged ???


Personal Relief 1,408


STATUTORIES


N.S.S.F 200


N.H.I.F 1,700


P.A.Y.E ???



1
Expert's answer
2022-02-02T13:11:18-0500

Gross pay=Basic pay + allowances

=163562+139675+10000

=313237


Taxable pay = 313237

from tax distribution table below:

Taxable income rate(for every 20 shilling)

1-75480 2

75430-150960 3

150961-226440 4

226441-301920 5

301921-377400 7


first bracket- 2/20*75480= 7548

second bracket- 3/20*75480=11322

third bracket - 4/20*75480= 15096

fourth bracket- 5/20*75480= 18870

fifth bracket-(313237-301920= 11317

7/20*11317= 3690.95


Tax charged= 7548+11322+15096+18870+3960.95

= 56796.96- 1408(personal relief)

= 55388.95


Net salary =Basic salary +(allowances-deductions)

= 313237-(55388.95+200+16356.2+200+1700

=313237-73845.15

=239391.85